# Hamiltonian Descent Algorithms for Optimization: Accelerated Rates via Randomized Integration Time

### v1

@@ C.2 Review of sampling algorithms
[…]
@@
[…]
@@
[…]
@@ Appendix D Hamiltonian flow with short-time integration
Given a continuous function $y(t)$ and positive scalars $m,\,L$ such that $0\leq y(t)\leq m+L\int_{0}^{t}(t-s)y(s)\,\mathrm{d}s$, we have $y(t)\leq m\cosh(\sqrt{L}t)$.
@@
$\|V_{t}\|\leq h\|\nabla f(X_{0})\|\cosh(\sqrt{L}t).$
@@ D.1 Proof of Theorem 1
D.2 Implementation oh HF-opt via leapfrog integrator as GD
@@ F.1 Proof of Theorem 2 (convergence of RHF under strong convexity)
Let ~$\widetilde{\mathcal{E}}_{t}=\mathbb{E}\left[f(X_{t})-f(x^{*})+\frac{\alpha}{10}\left\|X_{t}-x^{*}+\sqrt{\frac{5}{\alpha}}Y_{t}\right\|^{2}\right]$. Applying the chain rule, we have
˙$\dot{\mathcal{E}}_{t}=\sqrt{\frac{\alpha}{5}}\exp\left(\sqrt{\frac{\alpha}{5}}t\right)\widetilde{\mathcal{E}}_{t}+\exp\left(\sqrt{\frac{\alpha}{5}}t\right)\dot{\widetilde{\mathcal{E}}}_{t}.$
Thus, in order to show ˙$\dot{\mathcal{E}}_{t}\leq 0$, it suffices to show ~˙$\dot{\widetilde{\mathcal{E}}}_{t}\leq-\sqrt{\frac{\alpha}{5}}\widetilde{\mathcal{E}}_{t}$. Applying Lemma 13 with $\phi_{t}=\widetilde{\mathcal{E}}_{t}$, we obtain
~˙$\dot{\widetilde{\mathcal{E}}}_{t}$ $=\mathbb{E}\left[\langle\nabla f(X_{t}),Y_{t}\rangle+\left\langle\left(\begin{array}[]{c}\frac{\alpha}{5}\left(X_{t}-x^{*}+\sqrt{\frac{5}{\alpha}}Y_{t}\right)\\ \sqrt{\frac{\alpha}{5}}\left(X_{t}-x^{*}+\sqrt{\frac{5}{\alpha}}Y_{t}\right)\end{array}\right),\left(\begin{array}[]{c}Y_{t}\\ -\nabla f(X_{t})\end{array}\right)\right\rangle\right]$
@@ F.2 Proof of Theorem 3 (convergence of RHF under weak convexity)
We will show ˙$\dot{\mathcal{E}}_{t}\leq 0$. Applying Lemma 13 with $\phi_{t}=\mathcal{E}_{t}$, we obtain
˙$\dot{\mathcal{E}}_{t}$ $=\mathbb{E}\left[\frac{(t+1)^{2}}{4}\langle\nabla f(X_{t}),Y_{t}\rangle+\frac{t+1}{2}(f(X_{t})-f(x^{*}))\right]$
@@ G.1.1 For strongly convex functions
$=-L_{k+1}+\frac{\alpha^{\nicefrac{{3}}{{2}}}h}{432}\|x^{*}-x_{k+1}\|^{2}+\left(\frac{\alpha h}{36}-\frac{\alpha^{\nicefrac{{3}}{{2}}}h^{2}}{36}\right)\langle x_{k+1}-x^{*},\tilde{y}_{k+1}\rangle+\left(\frac{\sqrt{\alpha}h}{12}-\frac{\alpha h^{2}}{12}\right)\|\tilde{y}_{k+1}\|^{2}$
@@ G.2 Approximation error
Let $\hat{x}_{k+1}=\mathrm{Prox}_{h^{2}f}(x_{k+\frac{1}{2}})$ and $x_{k+1}=x_{k+\frac{1}{2}}-h^{2}\nabla f(x_{k+\frac{1}{2}})$. If $f$ is $L$-smooth and $h\leq\frac{1}{4^{\nicefrac{{1}}{{8}}}\cdot\sqrt{L}}$, then
$\|x_{k+1}-\hat{x}_{k+1}\|^{2}\leq 4L^{2}h^{8}\|\nabla f(x_{k+1})\|^{2}.$
@@
If $h\leq\frac{1}{4^{\nicefrac{{1}}{{8}}}\sqrt{L}}$, then rearranging the relation above yields
$\|x_{k+1}-\hat{x}_{k+1}\|^{2}\leq\frac{2L^{2}h^{8}}{1-2L^{4}h^{8}}\|\nabla f(x_{k+1})\|^{2}\leq 4L^{2}h^{8}\|\nabla f(x_{k+1})\|^{2}.$
@@ G.3.1 Proof of Theorem 4 (convergence of RHGD under strong convexity)
$\mathsf{II}$ $=\frac{\alpha h}{36}\left\langle\tilde{y}_{k+1}+\mathcal{G}_{k+1}^{h}-\frac{6}{\sqrt{\alpha}}\nabla f(x_{k+1}),x_{k+1}-x^{*}+\frac{6}{\sqrt{\alpha}}\tilde{y}_{k+1}\right\rangle$
@@
$\qquad+\frac{\alpha h}{36}\langle\mathcal{G}_{k+1}^{h},x_{k+1}-x^{*}\rangle+\frac{\sqrt{\alpha}h}{6}\langle\mathcal{G}_{k+1}^{h},\tilde{y}_{k+1}\rangle.$ (29)
@@
$\mathsf{III}$ $=-\frac{\alpha h^{2}}{72}\|\tilde{y}_{k+1}+\mathcal{G}_{k+1}^{h}-\frac{6}{\sqrt{\alpha}}\nabla f(x_{k+1})\|^{2}$
@@
$\qquad-\frac{\alpha h^{2}}{36}\left\langle\tilde{y}_{k+1}-\frac{6}{\sqrt{\alpha}}\nabla f(x_{k+1}),\mathcal{G}_{k+1}^{h}\right\rangle-\frac{\alpha h^{2}}{72}\|\mathcal{G}_{k+1}^{h}\|^{2}$
@@
$\qquad-\frac{\alpha h^{2}}{36}\langle\tilde{y}_{k+1},\mathcal{G}_{k+1}^{h}\rangle+\frac{\sqrt{\alpha}h^{2}}{6}\langle\nabla f(x_{k+1}),\mathcal{G}_{k+1}^{h}\rangle-\frac{\alpha h^{2}}{72}\|\mathcal{G}_{k+1}^{h}\|^{2}.$ (31)
@@
$\qquad+\frac{\sqrt{\alpha}h}{6}\langle\tilde{y}_{k+1},\mathcal{G}_{k+1}^{h}\rangle-\alpha h\langle\tilde{y}_{k+1},\mathcal{G}_{k+1}^{h}\rangle-\frac{\alpha h^{2}}{36}\langle\tilde{y}_{k+1},\mathcal{G}_{k+1}^{h}\rangle$ (33)
$\qquad+\frac{\alpha h}{36}\langle\mathcal{G}_{k+1}^{h},x_{k+1}-x^{*}\rangle-\left(\frac{\alpha}{2}+\frac{\alpha h^{2}}{72}\right)\|\mathcal{G}_{k+1}^{h}\|^{2}.$ (34)
@@
$\leq\frac{h^{2}}{8}\|\nabla f(x_{k+1})\|^{2}+\frac{2}{h^{2}}\|\mathcal{G}_{k+1}^{h}\|^{2}+\frac{h^{2}}{12}\|\nabla f(x_{k+1})\|^{2}+\frac{\alpha h^{2}}{12}\|\mathcal{G}_{k+1}^{h}\|^{2}$
$=\frac{5h^{2}}{24}\|\nabla f(x_{k+1})\|^{2}+\left(\frac{2}{h^{2}}+\frac{\alpha h^{2}}{12}\right)\|\mathcal{G}_{k+1}^{h}\|^{2}.$
@@
$\leq\frac{\sqrt{\alpha}h}{6}\|\tilde{y}_{k+1}\|^{2}+\frac{\sqrt{\alpha}h}{24}\|\mathcal{G}_{k+1}^{h}\|^{2}+\frac{\alpha h^{2}}{4}\|\tilde{y}_{k+1}\|^{2}+\alpha\|\mathcal{G}_{k+1}^{h}\|^{2}+\frac{\alpha h^{2}}{72}\|\tilde{y}_{k+1}\|^{2}+\frac{\alpha h^{2}}{72}\|\mathcal{G}_{k+1}^{h}\|^{2}$
$=\left(\frac{\sqrt{\alpha}h}{6}+\frac{19\alpha h^{2}}{72}\right)\|\tilde{y}_{k+1}\|^{2}+\left(\frac{\sqrt{\alpha}h}{24}+\alpha+\frac{\alpha h^{2}}{72}\right)\|\mathcal{G}_{k+1}^{h}\|^{2}.$
[…]
$\leq\frac{\sqrt{\alpha}}{72}\|\mathcal{G}_{k+1}^{h}\|^{2}+\frac{\alpha^{\nicefrac{{3}}{{2}}}h^{2}}{72}\|x^{*}-x_{k+1}\|^{2}-\left(\frac{\alpha}{2}+\frac{\alpha h^{2}}{72}\right)\|\mathcal{G}_{k+1}^{h}\|^{2}.$
Thus we have
[…]
$\qquad+\left(\frac{2}{h^{2}}+\frac{\alpha h^{2}}{12}+\frac{\sqrt{\alpha}h}{24}+\frac{\alpha}{2}+\frac{\sqrt{\alpha}}{72}\right)\|\mathcal{G}_{k+1}^{h}\|^{2}$
[…]
$\mathcal{L}_{k+1}-\mathcal{L}_{k}$ $\leq-\frac{\sqrt{\alpha}h}{6}\mathcal{L}_{k+1}-\frac{7\alpha^{\nicefrac{{3}}{{2}}}h^{2}}{432}\|x^{*}-x_{k+1}\|^{2}-\frac{17\alpha h^{2}}{72}\|\tilde{y}_{k+1}\|^{2}-\frac{5h^{2}}{24}\|\nabla f(x_{k+1})\|^{2}$
$\qquad+\left(\frac{2}{h^{2}}+\frac{\alpha h^{2}}{12}+\frac{\sqrt{\alpha}h}{24}+\frac{\alpha}{2}+\frac{\sqrt{\alpha}}{72}\right)\|\mathcal{G}_{k+1}^{h}\|^{2}$
$\leq-\frac{\sqrt{\alpha}h}{6}\mathcal{L}_{k+1}+\left(\frac{2}{h^{2}}+\frac{\alpha h^{2}}{12}+\frac{\sqrt{\alpha}h}{24}+\frac{\alpha}{2}+\frac{\sqrt{\alpha}}{72}\right)\|\mathcal{G}_{k+1}^{h}\|^{2}-\frac{5h^{2}}{24}\|\nabla f(x_{k+1})\|^{2}.$ (35)
@@
$\|\mathcal{G}_{k+1}^{h}\|^{2}\leq 2\cdot\left(1+\frac{1}{2}\right)\|x_{k+1}-\hat{x}_{k+1}\|^{2}\leq 3\|x_{k+1}-\hat{x}_{k+1}\|^{2}\leq 12L^{2}h^{8}\|\nabla f(x_{k+1})\|^{2}.$
Since we assume $L\geq 1$ and $\alpha\leq 1$, we have $h\leq\frac{1}{4^{\nicefrac{{1}}{{12}}}\sqrt{L}}\leq 1$, which implies
$\left(\frac{2}{h^{2}}+\frac{\alpha h^{2}}{12}+\frac{\sqrt{\alpha}h}{24}+\frac{\alpha}{2}+\frac{\sqrt{\alpha}}{72}\right)$ $\leq\left(2+\frac{1}{12}+\frac{1}{24}+\frac{1}{2}+\frac{1}{72}\right)\cdot\max\left(\frac{1}{h^{2}},\alpha h^{2},\sqrt{\alpha}h,\alpha,\sqrt{\alpha}\right)=\frac{95}{36h^{2}}.$
If we choose $h\leq\frac{1}{4\sqrt{L}}$, we have $\frac{95L^{2}h^{4}}{3}\leq\frac{95}{768}\leq\frac{5}{24}.$ Thus we obtain from (35):
$\mathcal{L}_{k+1}-\mathcal{L}_{k}$ $\leq-\frac{\sqrt{\alpha}h}{6}\mathcal{L}_{k+1}-h^{2}\left(\frac{5}{24}-\frac{95L^{2}h^{4}}{3}\right)\|\nabla f(x_{k+1})\|^{2}\leq-\frac{\sqrt{\alpha}h}{6}\mathcal{L}_{k+1},$

### v2

@@ C.2 Review of sampling algorithms
$\left\{\begin{aligned} y_{k}\mid x_{k}&\sim\tilde{\nu}^{Y\mid X}(\cdot\mid x_{k})=\mathcal{N}(x_{k},\eta I),\\ x_{k+1}\mid y_{k}&\sim\tilde{\nu}^{X\mid Y}(\cdot\mid y_{k}),\end{aligned}\right$. (Prox-S)
@@
$\left\{\begin{aligned} \mathrm{d}X_{t}&=Y_{t}\,\mathrm{d}t,\\ \mathrm{d}Y_{t}&=-\beta Y_{t}\,\mathrm{d}t-\nabla f(X_{t})\,\mathrm{d}t+\sqrt{2\beta}\,\mathrm{d}\mathrm{B}_{t},\end{aligned}\right$. (ULD)
@@
$\left\{\begin{aligned} \dot{X}_{t}&=Y_{t},\\ \dot{Y}_{t}&=-\beta Y_{t}-\nabla f(X_{t})-\nabla\log\mu_{t}(X_{t}),\end{aligned}\right$. (AIG)
@@ Appendix D Hamiltonian flow with short-time integration
Given a continuous function $y(t)$ and positive scalars m,Lm,\,L such that $0\leq y(t)\leq m+L\int_{0}^{t}(t-s)y(s)\,\mathrm{d}s$, we have $y(t)\leq m\cosh(\sqrt{L}t)$.
@@
$\|Y_{t}\|\leq h\|\nabla f(X_{0})\|\cosh(\sqrt{L}t).$
@@ D.1 Proof of Theorem 1
D.2 Implementation of HF-opt via leapfrog integrator as GD
@@ F.1 Proof of Theorem 2 (convergence of RHF under strong convexity)
Let $\widetilde{\mathcal{E}}_{t}=\mathbb{E}\left[f(X_{t})-f(x^{*})+\frac{\alpha}{10}\left\|X_{t}-x^{*}+\sqrt{\frac{5}{\alpha}}Y_{t}\right\|^{2}\right]$. Applying the chain rule, we have
$\dot{\mathcal{E}}_{t}=\sqrt{\frac{\alpha}{5}}\exp\left(\sqrt{\frac{\alpha}{5}}t\right)\widetilde{\mathcal{E}}_{t}+\exp\left(\sqrt{\frac{\alpha}{5}}t\right)\dot{\widetilde{\mathcal{E}}}_{t}.$
Thus, in order to show $\dot{\mathcal{E}}_{t}\leq 0$, it suffices to show $\dot{\widetilde{\mathcal{E}}}_{t}\leq-\sqrt{\frac{\alpha}{5}}\widetilde{\mathcal{E}}_{t}$. Applying Lemma 13 with $\phi_{t}=\widetilde{\mathcal{E}}_{t}$, we obtain
$\dot{\widetilde{\mathcal{E}}}_{t}$ $=\mathbb{E}\left[\langle\nabla f(X_{t}),Y_{t}\rangle+\left\langle\left(\begin{array}[]{c}\frac{\alpha}{5}\left(X_{t}-x^{*}+\sqrt{\frac{5}{\alpha}}Y_{t}\right)\\ \sqrt{\frac{\alpha}{5}}\left(X_{t}-x^{*}+\sqrt{\frac{5}{\alpha}}Y_{t}\right)\end{array}\right),\left(\begin{array}[]{c}Y_{t}\\ -\nabla f(X_{t})\end{array}\right)\right\rangle\right]$
@@ F.2 Proof of Theorem 3 (convergence of RHF under weak convexity)
We will show $\dot{\mathcal{E}}_{t}\leq 0$. Applying Lemma 13 with $\phi_{t}=\mathcal{E}_{t}$, we obtain
$\dot{\mathcal{E}}_{t}$ $=\mathbb{E}\left[\frac{(t+1)^{2}}{4}\langle\nabla f(X_{t}),Y_{t}\rangle+\frac{t+1}{2}(f(X_{t})-f(x^{*}))\right]$
@@ G.1.1 For strongly convex functions
$=-L_{k+1}+\frac{\alpha}{72}\|x^{*}-x_{k+1}\|^{2}+\left(\frac{\sqrt{\alpha}}{6}-\frac{\alpha h}{6}\right)\langle x_{k+1}-x^{*},\tilde{y}_{k+1}\rangle+\left(\frac{1}{2}-\frac{\sqrt{\alpha}h}{2}\right)\|\tilde{y}_{k+1}\|^{2}$
@@ G.2 Approximation error
Let $\hat{x}_{k+1}=\mathrm{Prox}_{h^{2}f}(x_{k+\frac{1}{2}})$ and $x_{k+1}=x_{k+\frac{1}{2}}-h^{2}\nabla f(x_{k+\frac{1}{2}})$. If $f$ is $L$-smooth and $h<\frac{1}{2^{\nicefrac{{1}}{{8}}}\cdot\sqrt{L}}$, then
$\|x_{k+1}-\hat{x}_{k+1}\|^{2}\leq\frac{2L^{2}h^{8}}{1-2L^{4}h^{8}}\|\nabla f(x_{k+1})\|^{2}.$
@@
If $h<\frac{1}{2^{\nicefrac{{1}}{{8}}}\sqrt{L}}$, then rearranging the relation above yields
$\|x_{k+1}-\hat{x}_{k+1}\|^{2}\leq\frac{2L^{2}h^{8}}{1-2L^{4}h^{8}}\|\nabla f(x_{k+1})\|^{2}.$
@@ G.3.1 Proof of Theorem 4 (convergence of RHGD under strong convexity)
$\mathsf{II}$ $=\frac{\alpha h}{36}\left\langle\tilde{y}_{k+1}+\frac{1}{h}\mathcal{G}_{k+1}^{h}-\frac{6}{\sqrt{\alpha}}\nabla f(x_{k+1}),x_{k+1}-x^{*}+\frac{6}{\sqrt{\alpha}}\tilde{y}_{k+1}\right\rangle$
@@
$\qquad+\frac{\alpha}{36}\langle\mathcal{G}_{k+1}^{h},x_{k+1}-x^{*}\rangle+\frac{\sqrt{\alpha}}{6}\langle\mathcal{G}_{k+1}^{h},\tilde{y}_{k+1}\rangle.$ (29)
@@
$\mathsf{III}$ $=-\frac{\alpha h^{2}}{72}\|\tilde{y}_{k+1}+\frac{1}{h}\mathcal{G}_{k+1}^{h}-\frac{6}{\sqrt{\alpha}}\nabla f(x_{k+1})\|^{2}$
@@
$\qquad-\frac{\alpha h^{2}}{36}\left\langle\tilde{y}_{k+1}-\frac{6}{\sqrt{\alpha}}\nabla f(x_{k+1}),\frac{1}{h}\mathcal{G}_{k+1}^{h}\right\rangle-\frac{\alpha}{72}\|\mathcal{G}_{k+1}^{h}\|^{2}$
@@
$\qquad-\frac{\alpha h}{36}\langle\tilde{y}_{k+1},\mathcal{G}_{k+1}^{h}\rangle+\frac{\sqrt{\alpha}h}{6}\langle\nabla f(x_{k+1}),\mathcal{G}_{k+1}^{h}\rangle-\frac{\alpha}{72}\|\mathcal{G}_{k+1}^{h}\|^{2}.$ (31)
@@
$\qquad+\frac{\sqrt{\alpha}}{6}\langle\tilde{y}_{k+1},\mathcal{G}_{k+1}^{h}\rangle-\alpha h\langle\tilde{y}_{k+1},\mathcal{G}_{k+1}^{h}\rangle-\frac{\alpha h}{36}\langle\tilde{y}_{k+1},\mathcal{G}_{k+1}^{h}\rangle$ (33)
$\qquad+\frac{\alpha}{36}\langle\mathcal{G}_{k+1}^{h},x_{k+1}-x^{*}\rangle-\frac{37\alpha}{72}\|\mathcal{G}_{k+1}^{h}\|^{2}.$ (34)
@@
(32) $=\langle\nabla f(x_{k+1}),\mathcal{G}_{k+1}^{h}\rangle+\frac{1}{6}\langle h\nabla f(x_{k+1}),\sqrt{\alpha}\mathcal{G}_{k+1}^{h}\rangle$
$\leq\frac{h^{2}}{8}\|\nabla f(x_{k+1})\|^{2}+\frac{2}{h^{2}}\|\mathcal{G}_{k+1}^{h}\|^{2}+\frac{h^{2}}{12}\|\nabla f(x_{k+1})\|^{2}+\frac{\alpha}{12}\|\mathcal{G}_{k+1}^{h}\|^{2}$
$=\frac{5h^{2}}{24}\|\nabla f(x_{k+1})\|^{2}+\left(\frac{2}{h^{2}}+\frac{\alpha}{12}\right)\|\mathcal{G}_{k+1}^{h}\|^{2}.$
(33) $=\frac{\sqrt{\alpha}}{6}\langle\tilde{y}_{k+1},\mathcal{G}_{k+1}^{h}\rangle+\alpha\left\langle(-h)\tilde{y}_{k+1},\mathcal{G}_{k+1}^{h}\right\rangle+\frac{\alpha}{36}\langle(-h)\tilde{y}_{k+1},\mathcal{G}_{k+1}^{h}\rangle$
$\leq\frac{\sqrt{\alpha}h}{6}\|\tilde{y}_{k+1}\|^{2}+\frac{\sqrt{\alpha}}{24h}\|\mathcal{G}_{k+1}^{h}\|^{2}+\frac{\alpha h^{2}}{4}\|\tilde{y}_{k+1}\|^{2}+\alpha\|\mathcal{G}_{k+1}^{h}\|^{2}+\frac{\alpha h^{2}}{72}\|\tilde{y}_{k+1}\|^{2}+\frac{\alpha}{72}\|\mathcal{G}_{k+1}^{h}\|^{2}$
$=\left(\frac{\sqrt{\alpha}h}{6}+\frac{19\alpha h^{2}}{72}\right)\|\tilde{y}_{k+1}\|^{2}+\left(\frac{\sqrt{\alpha}}{24h}+\alpha+\frac{\alpha}{72}\right)\|\mathcal{G}_{k+1}^{h}\|^{2}.$
(34) $=\frac{1}{36}\left\langle\alpha^{\nicefrac{{1}}{{4}}}h^{-1}\mathcal{G}_{k+1}^{h},\alpha^{\nicefrac{{3}}{{4}}}h(x_{k+1}-x^{*})\right\rangle-\frac{37\alpha}{72}\|\mathcal{G}_{k+1}^{h}\|^{2}$
$\leq\frac{\sqrt{\alpha}}{72h^{2}}\|\mathcal{G}_{k+1}^{h}\|^{2}+\frac{\alpha^{\nicefrac{{3}}{{2}}}h^{2}}{72}\|x^{*}-x_{k+1}\|^{2}-\frac{37\alpha}{72}\|\mathcal{G}_{k+1}^{h}\|^{2}.$
Since $\gamma_{k}\cdot h=\sqrt{\alpha}h\leq 1$ and $\alpha\leq 1$, we have
@@
$\qquad+\left(\frac{2}{h^{2}}+\frac{7\alpha}{12}+\frac{\sqrt{\alpha}}{24h}+\frac{\sqrt{\alpha}}{72h^{2}}\right)\|\mathcal{G}_{k+1}^{h}\|^{2}$
$=\frac{5h^{2}}{24}\|\nabla f(x_{k+1})\|^{2}+\left(\frac{\sqrt{\alpha}h}{6}+\frac{19\alpha h^{2}}{72}\right)\|\tilde{y}_{k+1}\|^{2}+\frac{\alpha^{\nicefrac{{3}}{{2}}}h^{2}}{72}\|x^{*}-x_{k+1}\|^{2}$
$\qquad+\frac{144+42\alpha h^{2}+3\sqrt{\alpha}h+\sqrt{\alpha}}{72h^{2}}\|\mathcal{G}_{k+1}^{h}\|^{2}$
$\leq\frac{5h^{2}}{24}\|\nabla f(x_{k+1})\|^{2}+\left(\frac{\sqrt{\alpha}h}{6}+\frac{19\alpha h^{2}}{72}\right)\|\tilde{y}_{k+1}\|^{2}+\frac{\alpha^{\nicefrac{{3}}{{2}}}h^{2}}{72}\|x^{*}-x_{k+1}\|^{2}+\frac{95}{36h^{2}}\|\mathcal{G}_{k+1}^{h}\|^{2}.$
@@
$\mathcal{L}_{k+1}-\mathcal{L}_{k}$ $\leq-\frac{\sqrt{\alpha}h}{6}\mathcal{L}_{k+1}-\frac{7\alpha^{\nicefrac{{3}}{{2}}}h^{2}}{432}\|x^{*}-x_{k+1}\|^{2}-\frac{17\alpha h^{2}}{72}\|\tilde{y}_{k+1}\|^{2}-\frac{5h^{2}}{24}\|\nabla f(x_{k+1})\|^{2}+\frac{95}{36h^{2}}\|\mathcal{G}_{k+1}^{h}\|^{2}$
$\leq-\frac{\sqrt{\alpha}h}{6}\mathcal{L}_{k+1}+\frac{95}{36h^{2}}\|\mathcal{G}_{k+1}^{h}\|^{2}-\frac{5h^{2}}{24}\|\nabla f(x_{k+1})\|^{2}.$ (35)
@@
$\|\mathcal{G}_{k+1}^{h}\|^{2}\leq 2\cdot\left(1+\frac{1}{2}\right)\|x_{k+1}-\hat{x}_{k+1}\|^{2}\leq 3\|x_{k+1}-\hat{x}_{k+1}\|^{2}\leq\frac{6L^{2}h^{8}}{1-2L^{4}h^{8}}\|\nabla f(x_{k+1})\|^{2}.$
If we choose $h\leq\frac{1}{4\sqrt{L}}$, we have $\frac{95L^{2}h^{4}}{6(1-2L^{4}h^{8})}\leq\frac{5}{24}.$ Thus we obtain from (35):
$\mathcal{L}_{k+1}-\mathcal{L}_{k}$ $\leq-\frac{\sqrt{\alpha}h}{6}\mathcal{L}_{k+1}-h^{2}\left(\frac{5}{24}-\frac{95L^{2}h^{4}}{6(1-2L^{4}h^{8})}\right)\|\nabla f(x_{k+1})\|^{2}\leq-\frac{\sqrt{\alpha}h}{6}\mathcal{L}_{k+1},$